40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 28 16:27:22 2012, ericq, Update, PSL, PBS and Half Wave plates introduced 
    Reply  Tue Aug 28 17:16:54 2012, ericq, Update, PSL, Power reduced! 
       Reply  Tue Aug 28 17:43:04 2012, Jenne, Update, PSL, Power reduced! 
       Reply  Wed Aug 29 17:02:41 2012, ericq, Update, PSL, Power reduced! 
          Reply  Wed Aug 29 17:35:34 2012, Koji, Update, PSL, Power reduced! 
    Reply  Tue Aug 28 17:51:39 2012, Jenne, Update, PSL, PBS and Half Wave plates introduced 
Message ID: 7308     Entry time: Wed Aug 29 17:02:41 2012     In reply to: 7297     Reply to this: 7310
Author: ericq 
Type: Update 
Category: PSL 
Subject: Power reduced! 

Quote:

We have now reduced the power being input to the MC from 1.25W to 10mW, and changed out the MC refl BS for a mirror. 

The power was reduced via the PBS we introduced in Entry 7295.

While we were in there, we took a look at the AS beam, which was looking clipped on the monitor. Jenne felt that it appears that the clipping seems to be occurring inside the vacuum, possibly on the faraday. This will be investigated during the vent. 

 The power has been increased to 20mW. We got the 10mW number from the linked elog entry above. However, after venting we were having problems locking the MC. Upon investigating past elog posts, we found that 20mW was actually the power used in the past. The MC will now autolock. 

ELOG V3.1.3-